OpenStudy (anonymous):

cos theta=-4/5 and 90 degrees< theta< 180. Find cos theta/2

OpenStudy (vishweshshrimali5):

First of all apply this: \[\large{\cos(2x) = 2\cos^2(x) - 1}\]

OpenStudy (anonymous):

I have no clue how to do that.

OpenStudy (vishweshshrimali5):

See, in place of x put theta/2

OpenStudy (anonymous):

How would you go from there?

OpenStudy (vishweshshrimali5):

First form the equation in which in place of x there is theta/2

OpenStudy (anonymous):

yeah i just did that. Cos(theta/2)= 2cos^2(theta/2) - 1

OpenStudy (vishweshshrimali5):

No no in left hand side, its 2x so 2*(theta/2) = theta

OpenStudy (anonymous):

so its theta=2cos^2(theta/2)-1

OpenStudy (vishweshshrimali5):

no its cos(theta)

OpenStudy (anonymous):

cos(theta)=2cos^2(theta/2)-1 Like this?

OpenStudy (vishweshshrimali5):

Yes

OpenStudy (vishweshshrimali5):

Now, you have cos(theta) so evaluate cos(theta/2)

OpenStudy (anonymous):

How would you do that?

OpenStudy (vishweshshrimali5):

See: \[\large{\cos(\theta) = 2\cos^2(\theta/2) - 1}\] \[\large{\implies 2cos^2(\theta/2) = cos(\theta)+1}\] \[\large{\implies cos^2(\theta/2) = (cos(\theta)+1)/2}\] \[\large{\implies cos(\theta/2) = \pm\sqrt{(cos(\theta)+1)/2}}\]

OpenStudy (vishweshshrimali5):

Now, you have cos(theta) , so calculate cos(theta/2).

OpenStudy (anonymous):

I dont know how to evaluate. I just started precalc and my teacher gave us a problem that would be on the test. if we get it we don't have to do it on the test.

OpenStudy (vishweshshrimali5):

Ok tell me can you add 1 to given value of cos(theta) ?

OpenStudy (anonymous):

Cos(theta)+1?

OpenStudy (vishweshshrimali5):

Yes what's the value of that ?

OpenStudy (anonymous):

is it 0?

OpenStudy (anonymous):

is the answer square root10 / 10